# CaMML: Context-Aware Multimodal Learner for Large Models

### v2

@@ 5.1 Quantitative Analyses
Here, we conducted an analysis of CaMML, using the CaMML-7B model for expeditious experiments, on the ScienceQA dataset. Our goal was to discern the significance of each module within the CaMML architecture, as well as the influence of critical hyperparameters such as the number of layers, the selection of $M$, and the quantity of retrieved context samples. The default settings for CaMML-7B baseline are as follows: the number of layers is set to 2 for all perceivers, $M=128$, $N=3$, and hidden size is set to 768.
@@ 5.1.2 Impact of Hyperparameters
We conducted a thorough exploration of the influence of hyperparameters on the performance of CaMML.
In Figure  3, we report the performance by varying the number of layers, hidden sizes, and query number $M$. Our observations reveal several key insights: (1) Increasing the number of layers in CaMML can have a detrimental effect on model performance. While we observe similar performance with both 2 layers and 12 layers, we opt for 2 layers due to its smaller model size; (2) Increasing the value of $M$ does not consistently lead to performance improvements; (3) Larger hidden sizes tend to yield more favorable results in our analysis.
Furthermore, we investigated the impact of the number of context samples, $N$, used in the training stage as well as the inference stage. We conduct experiments where we varied the value of $N$ during training and inference, and the corresponding performance is reported in Figure 4. We observe that: (1) It is easy for CaMML to accommodate a large number of shots; (2) Increasing the value of $N$ during inference does not consistently result in improved performance. Similar trends have been reported in previous works such as [3] and [55]. One plausible explanation for this phenomenon could be that a longer context might introduce complexity and potentially convolute the inference process; (3) The “mixed-shots” training strategy (Sec 4.2) has shown the potential to yield constant superior performance when compared to using a fixed value for $N$.
Figure 5: Visualization of context-aware CaMML vs. no-context LLaVA-1.5. Left: sketch drawing of the Great Wall. Right: depiction of the biology metamorphosis of a butterfly.
Figure 6: Visualization of CaMML vs. GPT-4V &\&& LLaVA-1.5. CaMML demonstrates a strong understanding of contexual sequences. We directly input contexual sequences (Left: consecutive Homer Simpson video frames and Right: Lionel Messi winning Ballon d’or in October 2023.) to CaMML perceiver without additional contexts.
@@ 5.2.1 The Importance of Context-Awareness
CaMML effectively processes a wide range of contextual inputs. In Figure 5, we compare the response from CaMML with context sample support and LLaVA-1.5, emphasizing the importance of context-awareness. The comparison illustrates that the inclusion of relevant context examples is crucial for ensuring the accuracy of answers. CaMML can not only infer using highly relevant context samples (The Great Wall example), but also draw insights from analogous domain-specific samples (The Metamorphosis example) and conduct multimodal analogy-based learning [5].
@@ 5.2.2 Handling Image Sequences
We also demonstrate that CaMML possesses the ability to handle image sequences, even in the absence of explicit training for this specific task. In Figure 6, instead of retrieval, we consider the last image of the image sequence as the $q$ and the preceding images as context samples. A comparison with LLaVA-1.5 and GPT-4V in Figure 6 on the QA task on image sequences reveals that while LLaVA-1.5 and GPT-4V struggles to fully comprehend the image sequences and learn from the priors, resulting in incorrect answers, CaMML can effectively understand the image sequences and give accurate responses. More importantly, CaMML can provide accurate inferences given access to the most up-to-date information regardless of LLM’s internal knowledge, yet both LLaVA-1.5 and GPT-4V lack the capability to incorporate real-time updates of world knowledge.
@@ 5.2.3 Tackling Multimodal Hallucination
Previous research indicates that RAG is effective in reducing hallucinations in content generation [48]. Here, we aim to investigate if CaMML can handle hallucinations in a multimodal setting. Figure 7 illustrates a multimodal QA task involving the recognition of similar objects: Muffins and Chihuahuas; Labradoodles and Fried Chicken. We treat the grid of images as a single image. It is evident that LLaVA-1.5 and GPT-4V face challenges in distinguishing between these objects, leading to hallucinated answers. In contrast, CaMML accurately answers the questions with the groundness of relevant context samples.
@@ A.1 Experimental Setup
ScienceQA [36] train split is used for ScienceQA finetuning and ablation study. In our instruction-tuning training, we adopt LLaVA-665K dataset, which contains LLaVA-158K, ShareGPT-40k [1], VQAv2 [4], GQA [18], OKVQA [39], OCRVQA [40], A-OKVQA [46], TextCaps [49], RefCOCO [58] and VG [24]. In our evaluation, VQAv2, GQA, TextVQA [50], MME [14], POPE [29], MM-Vet [59], ScienceQA, MMBench [34], MMBench-CN [34], SEED-Bench [27] and Vizwiz [16] are considered as benchmarks. We also conduct evaluation on COCO Caption [10], Flickr30k [41], OKVQA [39], A-OKVQA [46], and RefCOCO/+/g [58]. In our qualitative visualization, we adopt another 2M datastore as the source of context examples, which comprised from external resources incorporates 2,348K multimodal samples, ranging from captioning with BLIP-LAION’s 558K entries and Local Narrative [42], to knowledge-based QA with KVQA [45], narrative-driven QA from VCR [60] and Visual7W [68], visual grounding via RefCOCOPlus [58] and RefCOCOg [58], OCR from TextOCR [51], along with the in-domain LLaVA-665K set.
@@
We utilize Vicuna-7B/13B [65] as our foundation Large Language Model (LLM), ViT-L-14 architecture is used as the vision encoder. In detail, Vicuna-v1.3333https://huggingface.co/lmsys/vicuna-7b-v1.3, https://huggingface.co/lmsys/vicuna-13b-v1.3 and CLIP-ViT-L-14444https://huggingface.co/openai/clip-vit-large-patch14 [43] initialization is for ScienceQA finetuning and Ablation studies, Vicuna-v1.5555https://huggingface.co/lmsys/vicuna-7b-v1.5, https://huggingface.co/lmsys/vicuna-13b-v1.5, CLIP-ViT-L-14-336px [^6] https://huggingface.co/openai/clip-vit-large-patch14-336, and LLaVA-1.5 multimodal projector [^7] https://huggingface.co/liuhaotian/llava-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, https://huggingface.co/liuhaotian/llava-v1.5-mlp2x-336px-pretrain-vicuna-13b-v1.5 is initialized for instruction finetuning. For the source of contexts, the ImageBind [^8] https://dl.fbaipublicfiles.com/imagebind/imagebind_huge.pth [15]-Huge model is adopted to embed texts and images, computing their similarities for indexing top-k samples.
@@ A.2 CaMML Retrieval Setup
  • Utilize ImageBind [15] model to inference upon images and get corresponding visual embedding.
  • Utilize Faiss [21] to build datastore index for visual embedding, and bind each data source (image and text) to the index, ensuring the right one-to-one retrieval.
@@ A.3.2 Finegrained Evaluation
MMICL [64] FLANT5-XXL 67.9 49.2 71.6 73.0 66.7 57.2 77.2 - - - - - - -
mPLUG-Owl2 [57] LLaMA2-7B 66.5 32.2 72.4 60.9 68.6 60.1 79.4 59.5 28.8 64.8 48.7 60.1 50.3 76.0
SPHINX [30] LLaMA2-13B 67.2 33.1 67.3 58.3 74.4 59.4 80.7 58.6 21.2 61.8 43.5 62.1 58.7 73.6
LLaVA-1.5 [31]-7B Vicuna-7B 63.0 26.3 68.8 53.0 67.2 56.6 76.4 57.4 25.4 58.8 55.7 55.3 49.7 75.7
LLaVA-1.5 [31]-13B Vicuna-13B 68.2 44.1 67.3 60.0 72.0 59.4 82.1 61.9 36.4 65.8 49.6 62.1 59.4 75.0
@@ Appendix B Additional Cases
CaMML also highlights its strong localization capabilities, particularly in terms of reasoning the concept behind and high-level understanding. We utilize CaMML to detect objects by providing implicit descriptions and allowing CaMML to reason about the nature and location of these objects. As shown in Figure 10, CaMML is able to identify the objects with implicit descriptions (such as airplane: “The thing that carries hundreds of souls to the sky” or mirror: “The thing that makes two same dogs existing in this image”, etc..).

### v3

@@ 5.1 Quantitative Analyses
Here, we conducted an analysis of CaMML, using the CaMML-7B model for expeditious experiments, on the ScienceQA dataset. Our goal was to discern the significance of each module within the CaMML architecture, as well as the influence of critical hyperparameters such as the number of layers, the selection of $M$, and the quantity of retrieved context samples. The default settings for CaMML-7B baseline are as follows: the number of layers is set to 2 for all perceivers, $M=128$, $N=3$, and hidden size is set to 768. We also conduct experiments in Sec 5.1.3 about CaMML computation (inference speed & memory footprint) compared with baseline approach, where all tokens are directly fed into LLMs.
@@ 5.1.2 Impact of Hyperparameters
In Figure  3, we report the performance by varying the number of layers, hidden sizes, and query number $M$. Our observations reveal several key insights: (1) Increasing the number of layers in CaMML can have a detrimental effect on model performance. While we observe similar performance with both 2 layers and 12 layers, we opt for 2 layers due to its smaller model size; (2) Increasing the value of $M$ does not consistently lead to performance improvements; (3) Larger hidden sizes yield more favorable results in our analysis.
Furthermore, we investigated the impact of the number of context samples, $N$, used in the training stage as well as the inference stage. We conduct experiments where we varied the value of $N$ during training and inference, and the corresponding performance is reported in Figure 4. We observe that: (1) It is easy for CaMML to accommodate a large number of shots; (2) Increasing the value of $N$ during inference does not consistently result in improved performance. Similar trends have been reported in previous works such as Alayrac et al. (2022) and Yang et al. (2023). One plausible explanation for this phenomenon could be that a longer context might introduce complexity and potentially convolute the inference process; (3) The “mixed-shots" training strategy (Sec 4.2) has shown the potential to yield constant superior performance when compared to using a fixed value for $N$.
@@ 5.1.3 Inference Cost
Figure 5: Visualization of context-aware CaMML vs. no-context LLaVA-1.5. Left: sketch drawing of the Great Wall. Right: depiction of metamorphosis of a butterfly.
Figure 6: Visualization of CaMML vs. GPT-4V &\&& LLaVA-1.5. CaMML demonstrates a strong understanding of contexual sequences. We directly input contexual sequences (Left: consecutive Homer Simpson video frames and Right: Lionel Messi winning Ballon d’or in October 2023.) to CaMML w/o additional contexts.
@@ 5.2.1 The Importance of Context-Awareness
CaMML effectively processes a wide range of contextual inputs. In Figure 5, we compare the response from CaMML with context sample support and LLaVA-1.5, emphasizing the importance of context-awareness. The comparison illustrates that the inclusion of relevant context examples is crucial for ensuring the accuracy of answers. CaMML can not only infer using highly relevant context samples (The Great Wall example), but also draw insights from analogous domain-specific samples (The Metamorphosis example) and conduct multimodal analogy-based learning Antonietti (2012).
@@ 5.2.2 Handling Image Sequences
We also demonstrate that CaMML possesses the ability to handle image sequences, even in the absence of explicit training for this specific task. In Figure 6, instead of retrieval, we consider the last image of the image sequence as the $q$ and the preceding images as context samples. A comparison with LLaVA-1.5 and GPT-4V in Figure 6 on the QA task on image sequences reveals that while LLaVA-1.5 and GPT-4V struggle to fully comprehend the image sequences and learn from the priors, resulting in incorrect answers, CaMML can effectively understand the image sequences and give accurate responses. More importantly, CaMML can provide accurate inferences given access to the most up-to-date information regardless of LLM’s internal knowledge, yet both LLaVA-1.5 and GPT-4V lack the capability to incorporate real-time updates of world knowledge.
@@ 5.2.3 Tackling Multimodal Hallucination
Previous research indicates that RAG is effective in reducing hallucinations in content generation Shuster et al. (2021). Here, we aim to investigate if CaMML can handle hallucinations in a multimodal setting. Figure 7 illustrates a multimodal QA task involving the recognition of similar objects: Muffins and Chihuahuas; Labradoodles and Fried Chicken. We treat the grid of images as a single image. It is evident that LLaVA-1.5 and GPT-4V face challenges in distinguishing between these objects, leading to hallucinated answers. In contrast, CaMML accurately answers the questions with the groundness of relevant context samples.
@@ 6 Conclusion
7 Limitations
While CaMML effectively integrates multimodal retrieval into large models, the presence of irrelevant examples may impede its performance. In addition, CaMML maintains a datastore for experimental analysis, yet it’s unable to encompass all data domains, such as medical images. Consequently, when the data distribution in the datastore diverges significantly from user or test queries, CaMML struggles to leverage its retrieval-augmented capability effectively. Lastly, despite its proficiency in inference and memory utilization, CaMML demands substantial training on large language models (similar to LLaVA) and relies on access to a sizable datastore to demonstrate its effectiveness.
8 Potential Risks
CaMML entails certain potential risks, including: Environmental unfriendliness and escalated computational costs, particularly noticeable with larger models. Despite being entirely encapsulated within its pipeline, with users unable to access the datastore directly (limited to embedding and data entry index), there remains a possibility for users to utilize the CaMML I/O API to instruct LLM to generate texts that may raise concerns or cause privacy leakage.
@@ A.1 Experimental Setup
ScienceQA Lu et al. (2022a) train split is used for ScienceQA finetuning and ablation study. In our instruction-tuning training, we adopt LLaVA-665K dataset, which contains LLaVA-158K, ShareGPT-40k sha (2023), VQAv2 Antol et al. (2015), GQA Hudson and Manning (2019), OKVQA Marino et al. (2019), OCRVQA Mishra et al. (2019), A-OKVQA Schwenk et al. (2022), TextCaps Sidorov et al. (2020), RefCOCO Yu et al. (2016) and VG Krishna et al. (2017). In our evaluation, VQAv2, GQA, TextVQA Singh et al. (2019), MME Fu et al. (2023), POPE Li et al. (2023c), MM-Vet Yu et al. (2023), ScienceQA, MMBench Liu et al. (2023d), MMBench-CN Liu et al. (2023d), SEED-Bench Li et al. (2023a) and Vizwiz Gurari et al. (2018) are considered as benchmarks. We also conduct evaluation on COCO Caption Chen et al. (2015), Flickr30k Plummer et al. (2015), OKVQA Marino et al. (2019), A-OKVQA Schwenk et al. (2022), and RefCOCO/+/g Yu et al. (2016). In our qualitative visualization, we adopt another 2M datastore as the source of context examples, which comprised from external resources incorporates 2,348K multimodal samples, ranging from captioning with BLIP-LAION’s 558K entries and Local Narrative Pont-Tuset et al. (2020), to knowledge-based QA with KVQA Sanket Shah and Talukdar (2019), narrative-driven QA from VCR Zellers et al. (2019) and Visual7W Zhu et al. (2016), visual grounding via RefCOCOPlus Yu et al. (2016) and RefCOCOg Yu et al. (2016), OCR from TextOCR Singh et al. (2021), along with the in-domain LLaVA-665K set.
Summary of Evaluation Metrics.
We illustrate the evaluation metrics on Table 2:
  • Accuracy: VQAv2, GQA, Vizwiz, SQA-Image, TextVQA, MMBench, MMBench-CN, SEED.
  • GPT4-Assisted Evaluation Score: MM-Vet.
  • F1 Score (POPE Paper Sec 5.1 Metrics): POPE.
  • Figures in MME represent sum of the scores (which measures QA accuracy and accuracy+, illustrated in MME Paper Fu et al. (2023) Sec 2.2) of all MME perception subtasks, including existence, count, position, color, poster, celebrity, scene, landmark, artwork, and OCR. The full score of each subtask is 200, and that of all perception is 2000.
@@
We utilize Vicuna-7B/13B Zheng et al. (2023) as our foundation Large Language Model (LLM), ViT-L-14 architecture is used as the vision encoder. In detail, Vicuna-v1.3333https://huggingface.co/lmsys/vicuna-7b-v1.3, https://huggingface.co/lmsys/vicuna-13b-v1.3 and CLIP-ViT-L-14444https://huggingface.co/openai/clip-vit-large-patch14 Radford et al. (2021) initialization is for ScienceQA finetuning and Ablation studies, Vicuna-v1.5555https://huggingface.co/lmsys/vicuna-7b-v1.5, https://huggingface.co/lmsys/vicuna-13b-v1.5, CLIP-ViT-L-14-336px [^6] https://huggingface.co/openai/clip-vit-large-patch14-336, and LLaVA-1.5 multimodal projector [^7] https://huggingface.co/liuhaotian/llava-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, https://huggingface.co/liuhaotian/llava-v1.5-mlp2x-336px-pretrain-vicuna-13b-v1.5 is initialized for instruction finetuning. For the source of contexts, the ImageBind [^8] https://dl.fbaipublicfiles.com/imagebind/imagebind_huge.pth Girdhar et al. (2023)-Huge model is adopted to embed texts and images, computing their similarities for indexing top-k samples.
@@ A.2 CaMML Retrieval Setup
  • Utilize ImageBind Girdhar et al. (2023) model to inference upon images and get corresponding visual embedding.
  • Utilize Faiss Johnson et al. (2019b) to build datastore index for visual embedding, and bind each data source (image and text) to the index, ensuring the right one-to-one retrieval.
@@ A.3.2 Finegrained Evaluation
MMICL Zhao et al. (2023) FLANT5-XXL 67.9 49.2 71.6 73.0 66.7 57.2 77.2 - - - - - - -
mPLUG-Owl2 Ye et al. (2023) LLaMA2-7B 66.5 32.2 72.4 60.9 68.6 60.1 79.4 59.5 28.8 64.8 48.7 60.1 50.3 76.0
SPHINX Lin et al. (2023) LLaMA2-13B 67.2 33.1 67.3 58.3 74.4 59.4 80.7 58.6 21.2 61.8 43.5 62.1 58.7 73.6
LLaVA-1.5 Liu et al. (2023a)-7B Vicuna-7B 63.0 26.3 68.8 53.0 67.2 56.6 76.4 57.4 25.4 58.8 55.7 55.3 49.7 75.7
LLaVA-1.5 Liu et al. (2023a)-13B Vicuna-13B 68.2 44.1 67.3 60.0 72.0 59.4 82.1 61.9 36.4 65.8 49.6 62.1 59.4 75.0
@@ Appendix B Additional Cases
CaMML also highlights its strong localization capabilities, particularly in terms of reasoning the concept behind and high-level understanding. We utilize CaMML to detect objects by providing implicit descriptions and allowing CaMML to reason about the nature and location of these objects. As shown in Figure 10, CaMML is able to identify the objects with implicit descriptions (such as airplane: “The thing that carries hundreds of souls to the sky” or mirror: “The thing that makes two same dogs existing in this image”, etc.).